# Networked Communication for Mean-Field Games with Function Approximation and Empirical Mean-Field Estimation

### v1

@@ Appendix A Experiments
Experiments were conducted on a Linux-based machine with 2 x Intel Xeon Gold 6248 CPUs (40 physical cores, 80 threads total, 55 MiB L3 cache). We use the JAX framework to accelerate and vectorise our code. Random seeds are set in our code in a fixed way dependent on the trial number to allow easy replication of experiments - please see the code files in the Supplementary Materials for details.
A.1 Games
@@
We conduct numerical tests with four games (defined by the agents’ reward/transition functions), chosen for being particularly amenable to intuitive and visualisable understanding of whether the agents are learning behaviours that are appropriate and explainable for the respective objective functions. In all cases, rewards are normalised in [0,1] after they are computed.
@@
This is the inverse of the ‘exploration’ game in (Lauriere et al. 2022), where in our case agents are encouraged to gather together by the reward function $R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})=$ log$(\hat{\mu}_{t}(s^{i}_{t}))$. That is, agent $i$ receives a reward that is logarithmically proportional to the fraction of the population that is co-located with it at time $t$. We give the population no indication where they should cluster, agreeing this themselves over time.
@@
Define the magnitude of the horizontal and vertical distances between $x,y$ at $t$ as $dist^{h}_{t}(x,y)$ and $dist^{v}_{t}(x,y)$ respectively. The state $s^{i}_{t}$ now consists of agent $i$’s position $x^{i}_{t}$ and the ‘shark’s’ position $\phi_{t}$. At each time step, the shark steps towards the most populated grid point $x^{*}_{t}=\arg\max_{x\in\mathcal{S}}\hat{\mu}_{t}(x)$. A horizontal step is taken if $dist^{h}_{t}(\phi_{t},x^{*}_{t})$geq $dist^{v}_{t}(\phi_{t},x^{*}_{t})$, otherwise a vertical step is taken. Additionally, the shark moves in a random direction with probability 0.01. Agents are rewarded more for being further from the shark, and also for clustering with other agents. The reward function is given by
@@
Diffuse.
We conduct an additional experiment not mentioned in the main text, which is similar to the ‘exploration’ tasks in Lauriere et al. (2022), Wu et al. (2024) and other MFG works. In our version agents are rewarded for being located in more sparsely populated areas but only if they are stationary. The reward function is given by $R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})=r_{stationary}(-\hat{\mu}_{t}(s^{i}_{t}))$, where
@@ A.2 Experimental metrics
Exploitability
Works on MFGs most commonly use the exploitability metric to evaluate how close a given policy $\pi$ is to a NE policy $\pi^{*}$ (Lauriere et al. 2022; Perrin et al. 2020; Laurière et al. 2022; Algumaei et al. 2023; Pérolat et al. 2022b; Benjamin and Abate 2024; Wu et al. 2024). The metric usually assumes that all agents are following the same policy $\pi$, and quantifies how much an agent can benefit by deviating from $\pi$ by measuring the difference between the return given by $\pi$ and that of a $BR$ policy with respect to the distribution indicated by $\pi$:
Definition 9 (Exploitability of $\pi$)
@@
If $\pi$ has a large exploitability then an agent can significantly improve its return by deviating from $\pi$, meaning that $\pi$ is far from $\pi^{*}$, whereas an exploitability of 0 implies that $\pi=\pi^{*}$. Prior works conducting empirical testing have generally focused on the centralised setting, so this classical definition, as well as most evaluations, only consider exploitability when all agents are following a single policy $\pi_{k}$. However, Benjamin and Abate (2024) notes that purely independent agents, as well as networked agents, may have divergent policies $\pi^{i}_{k}\neq\pi^{j}_{k}$ $\forall i,k\in 1,\dots,N$, as in our own setting. We therefore are interested in the ‘exploitability’ of the population’s joint policy $\boldsymbol{\pi}$ := ($\pi^{1},\dots,\pi^{N}$) $\in\Pi^{N}$.
@@
Since we do not have access to the exact $BR$ policy as in some related works (Lauriere et al. 2022; Wu et al. 2024), we must instead approximate the exploitability, similarly to (Perrin et al. 2021; Benjamin and Abate 2024). We freeze the policy of all agents apart from a deviating agent, for which we store its current policy and then conduct 50 $k$ loops of policy improvement. To approximate the expectations in Def. 9, we take the best return of the deviating agent across 10 additional $k$ loops, as well as the mean of all the other agents’ returns across these same 10 loops. (While the policies of all non-deviating agents is $\pi_{k}$ in the centralised case, if the non-deviating agents do not share a single policy, then this method is in fact approximating the exploitability of their joint policy $\boldsymbol{\pi}^{-d}_{k}$, where $d$ is the deviating agent.) We then revert the agent back to its stored policy, before learning continues for all agents as per the main algorithm. Due to the expensive computations required for this metric, we evaluate it every second $k$ iteration of the main algorithm for Figs. 1, 2, 7, 8 and 9, and every fourth iteration for the population-dependent experiments.
@@
The exploitability metric has a number of limitations in our setting. Our approximation takes place via MOMD policy improvement steps (as in the main algorithm) for a independent, deviating agent while the policies of the rest of the population are frozen. As such, the quality of our approximation is limited by the number of policy improvement/expectation rounds, which must be restricted for the sake of running speed of the experiments. Moreover, since one of the findings of our paper is that networked agents can improve their policies faster than independent or centralised agents, especially when non-linear function approximation is used, it is arguably unsurprising that approximating the $BR$ by an independently deviating agent sometimes gives an unclear and noisy metric. This includes the exploitability going below zero, which should not be possible if the policies and distributions are computed exactly.
@@
Moreover, in coordination games (the setting on which we focus), agents benefit by following the same behaviour as others, and so a deviating agent generally stands to gain less from a $BR$ policy than it might in the non-coordination games on which many other works focus. For example, the return of a best-responding agent in the ‘push object’ game still depends on the extent to which other agents coordinate on which direction in which to push the box, meaning it cannot significantly increase its return by deviating. This means that the downward trajectory of the exploitability metric is less clear in our plots than in other works. This is likely why the approximated exploitability gets lower in the ‘disperse’ task in Fig. 8 than in the other tasks. Given the limitations presented by approximating exploitability, we also provide the second metric to indicate the progress of learning.
Average discounted return
@@
We record the average discounted return of the agents’ policies $\pi_{k}^{i}$ during the $M$ iterations - this allows us to observe that settings that converge to similar exploitability values may not have similar average agent returns, suggesting that some algorithms are better than others at finding not just NE, but preferable NE. See for example Figs. 1 and 7, where the networked agents converge to similar exploitability as the independent and centralised agents, but receive higher average returns.
A.3 Hyperparameters
@@
Gridsize 10x10 / 50x50 / 100x100 Experiments with population-dependent policies are run on the 10x10 grid (Figs. 3, 4, 5 and 6), while experiments on large state spaces are run on 50x50 and 100x100 grids (Figs. 1, 2, 7, 8 and 9).
Population 500 We chose 500 for our demonstrations to show that our algorithm can handle large populations, indeed often larger than those demonstrated in other mean-field works, especially for grid-world environments, while also being feasible to simulate wrt. time and computation constraints (Yongacoglu, Arslan, and Yüksel 2022; Cui and Koeppl 2021; Cui, Fabian, and Koeppl 2023; Guo et al. 2020; Subramanian and Mahajan 2019; Yang et al. 2018; Subramanian et al. 2020, 2022, 2021; Wu et al. 2024; Benjamin and Abate 2024).
@@
$C_{p}$ 1 As in Benjamin and Abate (2024), we choose this value to show the convergence benefits brought by even a single communication round, even in networks that may have limited connectivity; higher choices are likely to have even better performance.
@@
$\tau_{q}$ 0.03 We tested $\tau_{q}$ in {0.01,0.02,0.03,0.04,0.05}, as well as linearly decreasing $\tau_{q}$ from 0.05 $\rightarrow$ 0 as $k$ increases. However, only 0.03 gave stable increase in return. Note that this is the value also chosen in Vieillard, Pietquin, and Geist (2020).
@@
$cl$1 We use the same value as in Vieillard, Pietquin, and Geist (2020).
$\nu$ $L-1$ We tested $\nu$ in $\{1,4,20,L-1\}$. We found that in our setting, updating $\theta^{\prime}\leftarrow\theta$ once per $k$ iteration s.t. $\theta^{\prime}_{k+1,l}=\theta_{k,l}$ $\forall l$ gave sufficient learning that was similar to the other potential choices of $\nu$, so we do this for simplicity, rather than arbitrarily choosing a frequency to update $\theta^{\prime}$ during each $k$ loop. Setting the target to be the policy from the previous iteration is similar to the method in Lauriere et al. (2022). Whilst Wu et al. (2024) updates the target within the $L$ loops for stability, we do not find this to be a problem in our experiments.
Optimiser Adam As in Vieillard, Pietquin, and Geist (2020), we use the Adam optimiser with initial learning rate 0.01.
$\tau_{k}$ cf. comment $\tau_{k}$ increases linearly from 0.001 to 1 across the $K$ iterations. This is a simplification of the annealing scheme used in Benjamin and Abate (2024). Further optimising the annealing process may lead to better results.
@@
A.4 Additional experiments
Figure 5: ‘Push object’ task, population-dependent policies with global observability of mean-field distribution, 10x10 grid.
Figure 6: ‘Evade’ task, population-dependent policies with global observability of mean-field distribution, 10x10 grid.
Figure 7: ‘Target agreement’ task, population-independent policies, 100x100 grid.
@@ A.4 Additional experiments
We provide additional experiments on large grids in Figs. 7, 8 and 9. In the ‘target agreement’ task in Fig. 7, the networked agents all significantly outperform both centralised and independent agents in term of average return, despite the independent agents appearing to have slightly lower exploitability. This is because independent agents are hardly improving their policies at all, such that there is little a deviating agent can do to increase its return in this coordination game, such that exploitability appears low, despite this being an undesirable equilibrium (see Appx. A.2 for further discussion on the limited information provided by the exploitation metric). In the ‘cluster’ task in Fig. 9, the networked agents obtain significantly higher return than the independent agents. While centralised agents have the lowest exploitability, networked agents of almost all communication radii outperform them in terms of average return, indicating that the communication scheme helps populations reach better performing equilibria.
@@
In the additional ‘disperse’ task in Fig. 8, networked agents significantly outperform independent and centralised agents in terms of average return. They also outperform centralised agents in terms of exploitability, and significantly outperform independent agents in terms of exploitability.
Explanation of better performance by networked agents than centralised agents
By adopting policy parameters $\pi^{j}_{k+1}$ from neighbours with probability related to their estimated value $\sigma^{j}_{k+1}$, networked agents can adopt parameters that are estimated to be better performing than their own, which intuitively means networked agents improve their return quicker than agents learning entirely independently. Moreover, since the centralised case involves a single central agent training its Q-network and then pushing its update automatically to all the others, this single set of updated parameters will be taken on by all agents whether or not they are actually particularly well performing. Since the networked agents are more likely to select better performing parameters out of a variety of options according to our algorithm, we find networked agents usually in fact learn quicker than centralised agents, and indeed the benefit is even greater in the function approximation setting than in the tabular case given in Benjamin and Abate (2024). It is significant that our communication scheme not only allows us to avoid the undesirable assumption of a centralised learner, but even to outperform it.
Appendix B Related work
@@
MFGs are a quickly growing research area, so we only discuss the works most closely related to this present work, and instead refer the reader to Benjamin and Abate (2024) for detailed discussion around the setting of networked communication for MFGs, and to Laurière et al. (2022) for a broader survey of MFGs. Our work is most closely related to Benjamin and Abate (2024), which introduced networked communication to the infinite-horizon MFG setting. However, this work focuses only on tabular settings rather than using function approximation as in ours, and only addresses population-independent policies.
@@
Lauriere et al. (2022) uses Munchausen Online Mirror Descent (MOMD), similar to our method for learning with neural networks, but there are numerous differences to our setting: most relevantly, they study a finite-horizon episodic setting, where the mean-field distribution is updated in an exact way and an oracle supplies a centralised learner with rewards and transitions for it to learn a population-independent policy. Wu et al. (2024) uses MOMD to learn population-dependent policies, albeit also with a centralised method that exactly updates the mean-field distribution in a finite-horizon episodic setting. Perrin et al. (2022) learns population-dependent policies with function approximation in infinite-horizon settings like our own, but does so in a centralised, two-timescale manner without using the empirical mean-field distribution.
@@
Yongacoglu, Arslan, and Yüksel (2022) addresses decentralised learning from a continuous, non-episodic run of the empirical system using either full or compressed information about the mean-field distribution, but agents are assumed to receive this information directly, rather than estimating it locally as in the algorithm we now present. They also do not consider function approximation or inter-agent communication in their algorithms. In the closely related but distinct area of Mean-Field RL, Subramanian et al. (2020) does estimate the empirical mean-field distribution from the local neighbourhood, however agents are seeking to estimate the mean action rather than the mean-field distribution over states as in our MFG setting. Their agents also do not have access to a communication network by which they can improve their estimates.
Appendix C Extensions and future work

### v2

@@ Appendix C Experiments
Experiments were conducted on a Linux-based machine with 2 x Intel Xeon Gold 6248 CPUs (40 physical cores, 80 threads total, 55 MiB L3 cache). We use the JAX framework to accelerate and vectorise our code. Random seeds are set in our code in a fixed way dependent on the trial number to allow easy replication of experiments.
C.1 Games
@@
We conduct numerical tests with five games. All are defined by the agents’ reward/transition functions, and chosen for being particularly amenable to intuitive and visualisable understanding of whether the agents are learning behaviours that are appropriate and explainable for the respective objective functions. In all cases, rewards are normalised in [0,1] after they are computed.
@@
This is the inverse of the ‘exploration’ game in Laurière et al. [2022b], where in our case agents are encouraged to gather together by the reward function $R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})=$ log$(\hat{\mu}_{t}(s^{i}_{t}))$. That is, agent $i$ receives a reward that is logarithmically proportional to the fraction of the population that is co-located with it at time $t$. We give the population no indication where they should cluster, agreeing this themselves over time.
@@
Define the magnitude of the horizontal and vertical distances between $x,y$ at $t$ as $dist^{h}_{t}(x,y)$ and $dist^{v}_{t}(x,y)$ respectively. The state $s^{i}_{t}$ now consists of agent $i$’s position $x^{i}_{t}$ and a ‘shark’s’ position $\phi_{t}$. At each time step, the shark steps towards the most populated grid point according to the empirical mean-field distribution i.e. $x^{*}_{t}=\arg\max_{x\in\mathcal{S}}\hat{\mu}_{t}(x)$. A horizontal step is taken if $dist^{h}_{t}(\phi_{t},x^{*}_{t})$geq $dist^{v}_{t}(\phi_{t},x^{*}_{t})$, otherwise a vertical step is taken. As well as featuring a non-stationary distribution, we add ‘common noise’ to the environment, with the shark in a random direction with probability 0.01. Such noise that affects the local states of all agents in the same way, making the evolution of the distribution stochastic, makes population-independent policies sub-optimal Laurière et al. [2022a]. Agents are rewarded more for being further from the shark, and also for clustering with other agents. The reward function is given by
@@
Disperse.
This is similar to the ‘exploration’ tasks in Laurière et al. [2022b], Wu et al. [2024] and other MFG works. In our version agents are rewarded for being located in more sparsely populated areas but only if they are stationary. The reward function is given by $R(s^{i}_{t},a^{i}_{t},\hat{\mu}_{t})=r_{stationary}(-\hat{\mu}_{t}(s^{i}_{t}))$, where
@@ C.2 Experimental metrics
C.2.1 Exploitability
Works on MFGs most commonly use the exploitability metric to evaluate how close a given policy $\pi$ is to a NE policy $\pi^{*}$ Laurière et al. [2022b]; Perrin et al. [2020]; Laurière et al. [2022a]; Algumaei et al. [2023]; Pérolat et al. [2022]; Benjamin and Abate [2023]; Wu et al. [2024]. The metric usually assumes that all agents are following the same policy $\pi$, and quantifies how much an agent can benefit by deviating from $\pi$ by measuring the difference between the return given by $\pi$ and that of a $BR$ policy with respect to the distribution generated by $\pi$:
Definition 9 (Exploitability of $\pi$).
@@ C.2.1 Exploitability
If $\pi$ has a large exploitability then an agent can significantly improve its return by deviating from $\pi$, meaning that $\pi$ is far from $\pi^{*}$, whereas an exploitability of 0 implies that $\pi=\pi^{*}$. Prior works conducting empirical testing have generally focused on the centralised setting, so this classical definition, as well as most evaluations, only consider exploitability when all agents are following a single policy $\pi_{k}$. However, Benjamin and Abate [2023] notes that purely independent agents, as well as networked agents, may have divergent policies $\pi^{i}_{k}\neq\pi^{j}_{k}$ $\forall i,k\in 1,\dots,N$, as in our own setting. We therefore are interested in the ‘exploitability’ of the population’s joint policy $\boldsymbol{\pi}$ := ($\pi^{1},\dots,\pi^{N}$) $\in\Pi^{N}$.
@@
Since we do not have access to the exact $BR$ policy as in some related works Laurière et al. [2022b]; Wu et al. [2024], we must instead approximate the exploitability, similarly to Perrin et al. [2021]; Benjamin and Abate [2023]. We freeze the policy of all agents apart from a deviating agent, for which we store its current policy and then conduct 50 $k$ loops of policy improvement. To approximate the expectations in Def. 9, we take the best return of the deviating agent across 10 additional $k$ loops, as well as the mean of all the other agents’ returns across these same 10 loops. (While the policies of all non-deviating agents is $\pi_{k}$ in the centralised case, if the non-deviating agents do not share a single policy, then this method is in fact approximating the exploitability of their joint policy $\boldsymbol{\pi}^{-d}_{k}$, where $d$ is the deviating agent.) We then revert the agent back to its stored policy, before learning continues for all agents as per the main algorithm. Due to the expensive computations required for this metric, we evaluate it every second $k$ iteration of the main algorithm for Figs. 1, 2, 3, 4 and 5, and every fourth iteration for the population-dependent experiments.
@@
The exploitability metric has a number of limitations in our setting. Our approximation takes place via MOMD policy improvement steps (as in the main algorithm) for an independent, deviating agent while the policies of the rest of the population are frozen. As such, the quality of our approximation is limited by the number of policy improvement/expectation rounds, which must be restricted for the sake of running speed of the experiments. Moreover, since one of the findings of our paper is that networked agents can improve their policies faster than independent or centralised agents, especially when non-linear function approximation is used, it is arguably unsurprising that approximating the $BR$ by an independently deviating agent sometimes gives an unclear and noisy metric. This includes the exploitability going below zero, which should not be possible if the policies and distributions are computed exactly.
@@
Moreover, in coordination games (the setting for all tasks apart from the ‘disperse’ game), agents benefit by following the same behaviour as others, and so a deviating agent generally stands to gain less from a $BR$ policy than it might in the non-coordination games on which many other works focus. For example, the return of a best-responding agent in the ‘push object’ game still depends on the extent to which other agents coordinate on which direction in which to push the box, meaning it cannot significantly increase its return by deviating. This means that the downward trajectory of the exploitability metric is less clear in our plots than in other works. This is likely why the approximated exploitability gets lower in the non-coordination ‘disperse’ task in Fig. 5 than in the other tasks. Given the limitations presented by approximating exploitability, we also provide the second metric to indicate the progress of learning.
C.2.2 Average discounted return
@@
We record the average discounted return of the agents’ policies $\pi_{k}^{i}$ during the $M$ iterations - this allows us to observe that settings that converge to similar exploitability values may not have similar average agent returns, suggesting that some algorithms are better than others at finding not just NE, but preferable NE. See for example Figs. 3 and 1, where the networked agents converge to similar exploitability as the independent and centralised agents, but receive higher average returns.
C.3 Hyperparameters
@@
Gridsize 10x10 / 50x50 / 100x100 Experiments with population-dependent policies are run on the 10x10 grid (Figs. 6, 7, 8 and 9), while experiments on large state spaces are run on 50x50 and 100x100 grids (Figs. 1, 2, 3, 4 and 5).
Population 500 We chose 500 for our demonstrations to show that our algorithm can handle large populations, indeed often larger than those demonstrated in other mean-field works, especially for grid-world environments, while also being feasible to simulate wrt. time and computation constraints Yongacoglu et al. [2022]; Cui and Koeppl [2021]; Cui et al. [2023]; Guo et al. [2023]; Subramanian and Mahajan [2019]; Yang et al. [2018]; Ganapathi Subramanian et al. [2021, 2020]; Subramanian et al. [2022]; Wu et al. [2024]; Benjamin and Abate [2023].
@@
$C_{p}$ 1 As in Benjamin and Abate [2023], we choose this value to show the convergence benefits brought by even a single communication round, even in networks that may have limited connectivity; higher choices are likely to have even better performance.
@@
$\tau_{q}$ 0.03 We tested $\tau_{q}$ in {0.01,0.02,0.03,0.04,0.05}, as well as linearly decreasing $\tau_{q}$ from 0.05 $\rightarrow$ 0 as $k$ increases. However, only 0.03 gave stable increase in return. Note that this is the value also chosen in Vieillard et al. [2020].
@@
$cl$1 We use the same value as in Vieillard et al. [2020].
$\nu$ $L-1$ We tested $\nu$ in $\{1,4,20,L-1\}$. We found that in our setting, updating $\theta^{\prime}\leftarrow\theta$ once per $k$ iteration s.t. $\theta^{\prime}_{k+1,l}=\theta_{k,l}$ $\forall l$ gave sufficient learning that was similar to the other potential choices of $\nu$, so we do this for simplicity, rather than arbitrarily choosing a frequency to update $\theta^{\prime}$ during each $k$ loop. Setting the target to be the policy from the previous iteration is similar to the method in Laurière et al. [2022b]. Whilst Wu et al. [2024] updates the target within the $L$ loops for stability, we do not find this to be a problem in our experiments.
Optimiser Adam As in Vieillard et al. [2020], we use the Adam optimiser with initial learning rate 0.01.
$\tau^{comm}_{k}$ cf. comment $\tau^{comm}_{k}$ increases linearly from 0.001 to 1 across the $K$ iterations. This is a simplification of the annealing scheme used in Benjamin and Abate [2023]. Further optimising the annealing process may lead to better results.
@@
C.4 Additional experiments
C.4.1 Additional experiments on large grids
We provide additional experiments on large grids in Figs. 3, 4 and 5.
In the ‘target agreement’ task in Fig. 3, the networked agents generally have lower exploitability than both centralised and independent agents, and significantly outperform the other architectures in terms of average return. As before, the margin by which the networked agents can outperform the centralised agents is much greater than in Benjamin and Abate [2023], showing that the benefits of the communication scheme are even greater in non-tabular settings.
@@ C.4.1 Additional experiments on large grids
In the ‘cluster’ task in Fig. 4, the networked agents obtain significantly higher return than the independent agents. While centralised agents have the lowest exploitability, networked agents of almost all communication radii outperform them in terms of average return, indicating that the communication scheme helps populations reach better performing equilibria.
@@
In the ‘disperse’ task in Fig. 5, networked agents significantly outperform independent and centralised agents in terms of average return. They also outperform centralised agents in terms of exploitability, and significantly outperform independent agents in terms of exploitability. The fact that this happens in this non-coordination, competitive game shows that agents do have an incentive to communicate with each other even if they are self-interested.
C.4.2 Population-dependent policies in complex environments
Figure 8: ‘Push object’ task, population-dependent policies with global observability of mean field, 10x10 grid.
Figure 9: ‘Evade’ task, population-dependent policies with global observability of mean field, 10x10 grid.
We also showcase the ability of our algorithm to handle two more complex tasks, using population-dependent policies and estimated mean-field observations.
Figs. 6 and 7, where agents estimate the mean-field distribution via Alg. 3, differ minimally from Figs. 8 and 9, where agents directly receive the global mean-field distribution. This shows that our estimation algorithm allows agents to appropriately estimate the distribution, even with only one round of communication for agents to help each other improve their local counts. Only in the ‘push object’ task in Fig. 6, and there only with the smaller broadcast radii, do agents slightly underperform the returns of agents in the global observability case in Fig. 8, as is reasonable.
For the reasons given in Sup. Mat. C.2.1, the exploitability metric gives limited information in the ‘push object’ task in Fig. 6. In the ‘evade’ task in Fig. 7, exploitability suggests that centralised learners outperform the other cases. However, all of the networked cases significantly outperform the independent learners in terms of the average return to which they converge in both tasks. In the ‘push object’ task networked learners also significantly outperform centralised learners in all but the case with the smallest broad communication radius, while in the ‘evade’ task all networked cases perform similarly to the centralised case. Recall though that in the real world a centralised architecture is a strong assumption, a computational bottleneck and single point of failure.
Appendix D Related work
@@
MFGs are a quickly growing research area, so we only discuss the works most closely related to this present work, and instead refer the reader to Benjamin and Abate [2023] for detailed discussion around the setting of networked communication for MFGs, and to Laurière et al. [2022a] for a broader survey of MFGs. Our work is most closely related to Benjamin and Abate [2023], which introduced networked communication to the infinite-horizon MFG setting. However, this work focuses only on tabular settings rather than using function approximation as in ours, and only addresses population-independent policies.
@@
Laurière et al. [2022b] uses Munchausen Online Mirror Descent (MOMD), similar to our method for learning with neural networks, but there are numerous differences to our setting: most relevantly, they study a finite-horizon episodic setting, where the mean-field distribution is updated in an exact way and an oracle supplies a centralised learner with rewards and transitions for it to learn a population-independent policy. Wu et al. [2024] uses MOMD to learn population-dependent policies, albeit also with a centralised method that exactly updates the mean-field distribution in a finite-horizon episodic setting. Perrin et al. [2022] learns population-dependent policies with function approximation in infinite-horizon settings like our own, but does so in a centralised, two-timescale manner without using the empirical mean-field distribution. Zhang et al. [2024] also uses function approximation along a non-episodic path, but involves a generic central agent learning an estimate of the mean field rather than using an empirical population. Approaches that directly update an estimate of the mean field must be able to generate rewards from this arbitrary mean field, even if they otherwise claim to be oracle-free. They are thus inherently centralised algorithms and rely on strong assumptions that may not apply in real-world problems. Conversely, we are interested in practical convergence in online, deployed settings, where the reward is computed from the empirical finite population.
@@
Yongacoglu et al. [2022] addresses decentralised learning from a continuous, non-episodic run of the empirical system using either full or compressed information about the mean-field distribution, but agents are assumed to receive this information directly, rather than estimating it locally as in the algorithm we now present. They also do not consider function approximation or inter-agent communication in their algorithms. In the closely related but distinct area of Mean-Field RL, Ganapathi Subramanian et al. [2021] does estimate the empirical mean-field distribution from the local neighbourhood, however agents are seeking to estimate the mean action rather than the mean-field distribution over states as in our MFG setting. Their agents also do not have access to a communication network by which they can improve their estimates.
Appendix E Limitations and future work
Our work follows the gold standard in MFGs by presenting experiments on grid world toy environments, albeit we show our algorithms are able to handle much larger and more complex games than prior work. Nevertheless future work lies in moving from these environments to real-world settings. In Sec. 5 we give theoretical results showing that our networked algorithm can outperform a centralised alternative. We leave more general analysis, such proof of convergence and sample guarantees in the function approximation setting, for future work.